While the pressure remains constant at 689.5 kPa the volume of a system of air changes from 0.567 m3 to 0.283 m3. Determine the heat added/rejected


1
Expert's answer
2021-10-06T13:01:03-0400

"P=689.5 \; kPa \\

V_1= 0.567 \;m^3 \\

V_2 = 0.283 \;m^3"

Assume m=1

"Q_{sup} = ΔH = mc_pΔT \\

= 1.005 \;kJ/kgK(T_2-T_1)"

From ideal gas equation

"PV_1 =mRT_1 \\

T_1 = \frac{689.5 \times 0.567}{0.287 \times 1} = 1362.18 \;K \\

PV_2=mRT_2 \\

T_2 = \frac{689.5 \times 0.283}{0.287 \times 1} \\

T_2 = 679.89 \;K \\

Q_{sup} = ΔH = 1.005 \times (679.89-1362.18) = -685.70 \;kJ"